OpenStudy (anonymous):

How to factor -x^2-6x=0

OpenStudy (anonymous):

its a quadratic, do you know how to factor quadratics?

OpenStudy (anonymous):

lol what selman

OpenStudy (solomonzelman):

-x^2-6x=0 take out -1 -1(x^2+6x)=0 divide both sides by '-1' x^2+6x=0 take out x x(x+6)=0 so either x=0 or x+6=0 ---> x=-6 2 answers: 0 and -6